Hydrogen atoms in a sample are excited to n = 5 state and it is found that photons of all possible wavelengths are present in emission spectra. The minimum number of hydrogen atoms in the sample would be
 
a) infinite b) 6  c) 10 d) 5

Grade:12

1 Answers

Sanyam Chhoriya
5 years ago
n*(n-1)/2 = 5*(4-1)/2 = 10........................                                    
Hope u understand...............
 
Regards. ....Sanyam